The manufacturer of the ColorSmart-5000 television set claims that 95 percent of its sets last at least five years without needing a single repair. In order to test this claim, a consumer group randomly selects 396 consumers who have owned a ColorSmart-5000 television set for five years. Of these 396 consumers, 311 say that their ColorSmart-5000 television sets did not need repair, while 85 say that their ColorSmart-5000 television sets did need at least one repair.

(a) Letting p be the proportion of ColorSmart-5000 television sets that last five years without a single repair, set up the null and alternative hypotheses that the consumer group should use to attempt to show that the manufacturer’s claim is false.

H0 :p   (Click to select)   >   ≥   <   ≠  .95 versus Ha : p   (Click to select)   ≠   >   ≥   <  .95.

(b) Use critical values and the previously given sample information to test the hypotheses you set up in part a by setting α equal to .10, .05, .01, and .001. How much evidence is there that the manufacturer’s claim is false?

  (Click to select)   Reject   Do not reject  H0 at each value of α;     (Click to select)   strong   very strong   none   extremely strong   some  evidence.
Answer #1

n = 396

x = 311

p̄ = x/n = 0.7854

a)

Null and Alternative hypothesis:

Ho : p = 0.95

H1 : p ≠ 0.95

b)

Test statistic:

z = (p̄ -p)/√(p*(1-p)/n) = (0.7854 - 0.95)/√(0.95 * 0.05/396) = -15.0333

Critical value :

z(0.10) = ABS(NORM.S.INV(0.10/2) = 1.645

z(0.05) = ABS(NORM.S.INV(0.05/2) = 1.96

z(0.01) = ABS(NORM.S.INV(0.01/2) = 2.576

z(0.001) = ABS(NORM.S.INV(0.001/2) = 3.291

Reject H0 at each value of α; extremely strong